Calculate the B/C ratio for a government investment of $2,000 which gives $800 at the end of every year 5 year period. The interest rate is 10%.
Answer #1

B/C ratio = PV of benefits/PV of cost

= 800/(1+0.10)^1+800/(1+0.10)^2+800/(1+0.10)^3+800/(1+0.10)^4+800/(1+0.10)^5 / 2000

= 727.27+661.15+601.05+546.41+496.73 / 2000

= 3032.61 / 2000

= 1.51